OpenStudy (anonymous):

which expression when evaluted, yields a true statement? A. 14-2m line under < 8 + 3m when m=1 B. 5x - 1< 3x + 7 when x=4 C. 5X - X/3 line under > 15 when x=3 D. 2x^2-3x + 1 < 16x- x2/4 when x=2

OpenStudy (anonymous):

which expression when evaluted, yields a true statement? A. 14-2m line under < 8 + 3m when m=1 B. 5x - 1< 3x + 7 when x=4 C. 5X - X/3 line under > 15 when x=3 D. 2x^2-3x + 1 < 16x- x2/4 when x=2

OpenStudy (anonymous):

@phi

OpenStudy (phi):

if you want to use an equivalent to \( \le \), use <= (means "less than or equal")

OpenStudy (anonymous):

oh okay so answer would be...

OpenStudy (anonymous):

b?

OpenStudy (phi):

replace "m" with the number. Then simplify using arithmetic remember a <= b means a is smaller than b or equal to b and a>=b means a is bigger than b or equal to b (of course, you need a and b to be numbers to decide)

OpenStudy (anonymous):

so the answer would be b?

OpenStudy (anonymous):

@sunnyshores